# Distributed Computing From First Principles

### v1

@@ 4.2 Election Algorithm
In distributed systems, allowing multiple nodes to propose values concurrently can introduce complexities in coordinating these proposals and can lead to increased contention for system resources. To streamline the consensus process and minimize such conflicts, it is often advantageous to designate a single node as the leader. This leader then assumes the role of a central coordinator, responsible for initiating and managing the agreement process. Our approach to leadership election relies on an underlying failure detector. This is a crucial component that continuously monitors the health and responsiveness of the participating nodes to identify potential failures or unresponsiveness of the currently designated leader [4, 37]. (For a more detailed explanation of failure detectors, please refer to the preceding section.)
We are using a variation of the bully algorithm which is described below [37]. The algorithm does the following:
@@ 4.3 Raft Algorithm
The Paxos algorithm, while foundational for achieving consensus in distributed systems, is often perceived by software engineers as complex and challenging to reason about. This complexity motivated the invention of the Raft algorithm, designed to be a more understandable alternative while providing equivalent functionality. Raft is described as a leader election-based sequence Paxos, essentially integrating the concepts of Paxos, a replicated log, and a designated leader [4, 12]. The elected leader acts as the sole proposer, centralizing coordination and minimizing potential contention. To ensure fault tolerance, Raft employs leader election, which is triggered if the current leader is detected as dead.
@@
  • Can an election choose multiple leaders [12]? Yes, although Raft’s mechanisms are designed to strongly discourage this, transient network issues or timing-dependent failures could theoretically lead to split votes and the possibility of multiple nodes believing they are leaders for a short period.
  • Can an election fail to choose a leader [12]? Yes, scenarios such as all candidates failing or a persistent split vote across the cluster could result in no single candidate obtaining a majority, thus leading to a failed election.
@@ 4.4 Stabilization Algorithm
This algorithm permits the return of a distributed computation to a ’correct behavior’ from a perturbed state [33]. This perturbation may be due to failures, such as those affecting networks and nodes. Self-repairing algorithms guarantee correct behavior even in the presence of failures [45]. Robust algorithms, as discussed in Chapter 4, have a fixed bound on the number and types of failures they can tolerate while exhibiting graceful degradation. In contrast, stabilization algorithms allow for any number of failures, with an expected delay before incorrect operation may occur until the failure is repaired and the system returns to a correct state. Stabilization algorithms are ideal for transient errors, as discussed in Section 3.4. Most robust algorithms follow a crash-stop model, whereas stabilization algorithms follow a crash-recovery model. Stabilization can be achieved through several means, including eliminating illegal states (domain restriction) and snapshotting.
@@
Snapshotting involves saving the configuration states of a distributed computation for later retrieval [43, 45]. When an error is detected, the system can revert to a previously known good configuration saved in the snapshot (a change point). Snapshotting is desirable when manual intervention to resolve errors is costly. Retrieving the saved change point when an illegal state is reached can guarantee self-healing in the distributed computing system. For reliability, it is essential that the saved change point has not been compromised by a malicious node, ensuring the system’s return to correct behavior.
@@ 4.5 Byzantine Protocol (BFT)
The resiliency discussed in SubSections 4.1, 4.3 would fail if there is $<\frac{1}{2}$ of the number of live processes. More robust algorithms can handle cases of failure exceeding the tolerance bound of traditional Paxos and raft algorithms. Some algorithms have mechanisms to handle cases where the failed process can take arbitrary actions such as sending bad messages and activating unnecessary state transitions, and even denial of service can affect a distributed system. It does not prevent some interruption, but it can mitigate the attack and ensure that an agreement can still be reached.
These algorithms with robust error handling included in the consensus algorithm are known as Byzantine protocols. These allow for building distributed systems that function in chaotic channels with significant disruption, as we can achieve a quorum with at least $\frac{1}{3}$ of the number of live processes. We consider scenarios with parallel leaders and also scenarios with a single leader serving as the proposer. There are several papers describing BFT [18, 41] with varying tolerance guarantees.
@@ 4.6 Distributed Commit and Transaction
Let’s set up a simple, single-system banking scenario with two accounts (A and B) to prepare for our discussion of distributed protocols in SubSection 4.6.1. We will illustrate a basic transfer using a flat transaction (Figure 4.7) and a more complex breakdown using a nested transaction (Figure LABEL:nested_banking_transaction).
@@
Figure LABEL:nested_banking_transaction presents a visual representation of a banking fund transfer implemented as a nested transaction, where the main operation is divided into a series of interdependent sub-operations. The overall process begins with a ”Start Transaction” action, which in turn triggers the first sub-transaction, labeled ”Start Sub-Transaction: Verify Funds.” The purpose of this initial sub-process is to confirm the availability of adequate funds through a ”CheckBalance” operation. Based on the outcome of this check, the system either proceeds to a ”SufficientFunds” state, culminating in a successful ”Commit Sub-Transaction: Verify Funds,” or encounters ”InsufficientFunds,” leading to an ”Abort Sub-Transaction: Verify Funds.”. The subsequent step, ”Start Sub-Transaction: Transfer Funds,” is initiated solely upon the successful completion (commit) of the ”Verify Funds” sub-transaction. This second sub-process involves the core actions of ”Debit Account A” and ”Credit Account B.” A successful execution of both these actions results in a ”Commit Sub-Transaction: Transfer Funds.” Conversely, any failure during either the debit or credit operation leads to an ”Abort Sub-Transaction: Transfer Funds.”. The success of the entire transaction, spanning from the initial ”Start Transaction” to the final ”Commit Transaction,” is contingent upon the successful commitment of both constituent sub-transactions. If either the ”Abort Sub-Transaction: Verify Funds” or the ”Abort Sub-Transaction: Transfer Funds” occurs, the entire operation is rolled back, resulting in an ”Abort Transaction.” This mechanism ensures the integrity of the data by reverting any partial modifications. Consequently, the final status of the primary transaction is directly determined by the combined outcomes of its sub-transactions.
@@ 4.6.1 Atomic Commit Protocols
Two-Phase Commit (2PC) protocol permits any party (participant) can decide to roll back its transaction (due to failure) and the entire transaction is globally reverted [20, 43]. This requires a coordinator to facilitate concurrency control. The algorithm has two phases [20]: in the first phase, each party votes to commit or abort a transaction. Once, a party has voted to commit a transaction, it is binding and immutable. After voting each party goes to the ”prepared” state. In the second phase, each party in the transaction jointly executes the decision. If any party aborts or fails, then the overall transaction is aborted. The two phases ensure that all parties reach the same decision on committing or rolling back the transaction.
@@
Figure 4.9: Two-Phase Commit [20].
@@
Two-Phase Commit fails when the coordinator is dead and participants in the transaction cannot determine the state of the transaction resulting in longer polling intervals as parties try to connect and get information from the coordinator. Three-Phase Commit solves the problem by taking more rounds [20, 43] as we perform a precommit round and obtain ACKS before proceeding to commit.
@@
Figure 4.10: Three-Phase Commit [20].
@@ 5.1 Gossip Algorithm
The Gossip protocol [25, 29, 30, 32] presents a robust and widely applicable anti-entropy mechanism that can be effectively employed in distributed systems to promote data consistency [37]. This protocol leverages a probabilistic approach to disseminate information (messages) across a network, drawing a compelling analogy to the way rumors or information propagate through a social network. A significant advantage of the Gossip algorithm is its inherent fault tolerance and resilience to failures, even in the face of substantial and unpredictable network disruptions or node outages. While the epidemic approach, characterized by the rapid and potentially ubiquitous spread of information to all reachable nodes, is a common and intuitive implementation strategy, Gossip protocols can also be implemented using non-epidemic strategies that strategically leverage overlay network structures or maintain partial views of the system’s membership to optimize message dissemination and resource utilization.
@@ 5.2 CRDT
Conflict-free Replicated Data Types (CRDTs) represent another powerful and increasingly popular class of anti-entropy mechanisms in distributed systems. Riak, a distributed NoSQL database, was a notable pioneer in the widespread adoption and practical application of CRDTs [5, 37]. A CRDT is a specialized type of data structure meticulously designed to maintain a consistent state across multiple independent replicas, regardless of the order in which concurrent operations are executed on those individual replicas. This inherent order-independence property makes CRDTs exceptionally well-suited for scenarios involving collaborative applications and seamless remote synchronization (replication) of data across numerous geographically distributed devices, a task that can be inherently complex to manage for data consistency using traditional concurrency control mechanisms. Mathematically, CRDTs can be formally characterized as partially ordered monoids that possess a well-defined lattice structure. To ensure their conflict-free nature, they must satisfy the algebraic properties of commutativity (the order in which operations are applied does not affect the final state) and idempotence (applying the same operation multiple times has the same effect as applying it only once).
@@
A diverse range of CRDTs has been developed, each tailored to manage different types of data and support specific sets of operations. Common examples include map CRDTs (such as the Last-Write-Wins Map - LWW-Map), set CRDTs (such as the Grow-Only Set - G-Set), and many other specialized data structures designed for various consistency and concurrency control requirements [5, 37].
@@
@TODO: IMPLEMENTATIONS
@@ 5.5.2 Multiple Description Coding
111 See an illustrative example of optimizing a distributed data store for operation at significant scale in the Uber blog
@@
Peer-to-peer computing offers notable benefits, including reduced storage and computational costs, achieved by distributing the burden of computation and data storage across several participating nodes. The applications of peer-to-peer computing span a wide range of domains, including distributed file systems (e.g., OceanStore [31], PAST [24], CFS [21]), file sharing networks (e.g., BitTorrent [1], FastTrack [2]), and large-scale Grid computing platforms [3].
@@ 6.1 Resilient Overlay Networks
An overlay network provides a dynamic, just-in-time network infrastructure built on top of an existing network layer (like the Internet) to facilitate decentralized communication in a scalable manner. It achieves this by implementing its own routing and packet transfer functionalities at the application level, abstracting away the complexities of the underlying physical network [16, 22].
Overlay networks must be designed to handle a high rate of client disconnections, which can occur due to various factors such as poor network conditions, limited available bandwidth, and the need for dynamic balancing of processing and bandwidth capacity across the participating nodes [17]. Consequently, there is a significant need for the development of resilient overlay networks that enhance the reliability of the network and make it robust to node and link failures [22]. Nodes within resilient overlay networks often incorporate self-healing capabilities. Such nodes can perform quality of service (QoS) checks on the communication links they use and dynamically re-route packets as necessary to circumvent failures and reduce latency.
@@
When communicating across different Internet Service Providers (ISPs), one might naively assume that Border Gateway Protocol (BGP) routing records are consistently up-to-date and reflect the optimal paths. However, this is often not the case, as the entries in an ISP’s routing table can be influenced by a multitude of factors, including financial agreements between ISPs, the time it takes for BGP to converge after a routing update, and even political considerations [22].
The routing table maintained at the application level by the overlay network can potentially identify and utilize more efficient and reliable communication links compared to the underlying BGP routes. Content Delivery Networks (CDNs) are a prominent example of how overlay networks are leveraged to improve offsite caching and the efficient delivery of web resources to end-users [22].
@@ 6.2 Unstructured P2P Systems
Unstructured overlay networks are characterized by the absence of a predefined, deterministic structure for organizing nodes and data. Consequently, they do not inherently provide efficient query-matching mechanisms for locating specific data within the network [17]. Examples of P2P systems that utilize unstructured overlays with a random graph topology include Gnutella [44], the now-defunct Kazaa, and Freenet [19]. In such networks, a typical search query is flooded across the network to a subset or all neighboring nodes. Each node receiving the query then compares it against the metadata or content it holds.
@@ 6.3 Structured P2P Systems
In contrast to unstructured P2P systems, structured overlay networks impose a specific, well-defined structure on the network topology to facilitate efficient query matching and data discovery [17]. This efficient search capability is achieved by enforcing constraints on how the network is constructed and how data is indexed and located. In a structured P2P system, each node is typically associated with a unique key or identifier, and the actual data or metadata is stored at the node whose key is closest to the key of the data item. Searching for a specific data item involves querying the network using its associated key. The network’s underlying structure, often based on sophisticated data structures like Distributed Hash Tables (DHTs), ensures that queries can be routed efficiently to the node(s) likely to hold the desired information. Prominent examples of P2P systems that leverage structured overlay networks include Chord [42], Tapestry [49], Kademlia [34], and Pastry [39].

### v2

@@ 4.2 Election Algorithm
In distributed systems, allowing multiple nodes to propose values concurrently can introduce complexities in coordinating these proposals and can lead to increased contention for system resources. To streamline the consensus process and minimize such conflicts, it is often advantageous to designate a single node as the leader. This leader then assumes the role of a central coordinator, responsible for initiating and managing the agreement process. Our approach to leadership election relies on an underlying failure detector. This is a crucial component that continuously monitors the health and responsiveness of the participating nodes to identify potential failures or unresponsiveness of the currently designated leader [kthcourseds, petrov2019]. (For a more detailed explanation of failure detectors, please refer to the preceding section.)
We are using a variation of the bully algorithm which is described below [petrov2019]. The algorithm does the following:
@@ 4.3 Raft Algorithm
The Paxos algorithm, while foundational for achieving consensus in distributed systems, is often perceived by software engineers as complex and challenging to reason about. This complexity motivated the invention of the Raft algorithm, designed to be a more understandable alternative while providing equivalent functionality. Raft is described as a leader election-based sequence Paxos, essentially integrating the concepts of Paxos, a replicated log, and a designated leader [kthcourseds, alinush2015]. The elected leader acts as the sole proposer, centralizing coordination and minimizing potential contention. To ensure fault tolerance, Raft employs leader election, which is triggered if the current leader is detected as dead.
@@
  • Can an election choose multiple leaders [alinush2015]? Yes, although Raft’s mechanisms are designed to strongly discourage this, transient network issues or timing-dependent failures could theoretically lead to split votes and the possibility of multiple nodes believing they are leaders for a short period.
  • Can an election fail to choose a leader [alinush2015]? Yes, scenarios such as all candidates failing or a persistent split vote across the cluster could result in no single candidate obtaining a majority, thus leading to a failed election.
@@ 4.4 Stabilization Algorithm
This algorithm permits the return of a distributed computation to a ’correct behavior’ from a perturbed state [lynch1993]. This perturbation may be due to failures, such as those affecting networks and nodes. Self-repairing algorithms guarantee correct behavior even in the presence of failures [tel2000]. Robust algorithms, as discussed in Chapter 4, have a fixed bound on the number and types of failures they can tolerate while exhibiting graceful degradation. In contrast, stabilization algorithms allow for any number of failures, with an expected delay before incorrect operation may occur until the failure is repaired and the system returns to a correct state. Stabilization algorithms are ideal for transient errors, as discussed in Section 3.4. Most robust algorithms follow a crash-stop model, whereas stabilization algorithms follow a crash-recovery model. Stabilization can be achieved through several means, including eliminating illegal states (domain restriction) and snapshotting.
@@
Snapshotting involves saving the configuration states of a distributed computation for later retrieval [tanenbaum2007, tel2000]. When an error is detected, the system can revert to a previously known good configuration saved in the snapshot (a change point). Snapshotting is desirable when manual intervention to resolve errors is costly. Retrieving the saved change point when an illegal state is reached can guarantee self-healing in the distributed computing system. For reliability, it is essential that the saved change point has not been compromised by a malicious node, ensuring the system’s return to correct behavior.
@@ 4.5 Byzantine Protocol (BFT)
The resiliency discussed in SubSections 4.1, 4.3 would fail if there is $\frac{1}{2}$ of the number of live processes. More robust algorithms can handle cases of failure exceeding the tolerance bound of traditional Paxos and raft algorithms. Some algorithms have mechanisms to handle cases where the failed process can take arbitrary actions such as sending bad messages and activating unnecessary state transitions, and even denial of service can affect a distributed system. It does not prevent some interruption, but it can mitigate the attack and ensure that an agreement can still be reached.
These algorithms with robust error handling included in the consensus algorithm are known as Byzantine protocols. These allow for building distributed systems that function in chaotic channels with significant disruption, as we can achieve a quorum with at least $\frac{1}{3}$ of the number of live processes. We consider scenarios with parallel leaders and also scenarios with a single leader serving as the proposer. There are several papers describing BFT [CastroL02, Stathakopoulou2022] with varying tolerance guarantees.
@@ 4.6 Distributed Commit and Transaction
Let’s set up a simple, single-system banking scenario with two accounts (A and B) to prepare for our discussion of distributed protocols in SubSection 4.6.1. We will illustrate a basic transfer using a flat transaction (Figure 4.7) and a more complex breakdown using a nested transaction (Figure 4.8).
@@
Figure 4.8 presents a visual representation of a banking fund transfer implemented as a nested transaction, where the main operation is divided into a series of interdependent sub-operations. The overall process begins with a ”Start Transaction” action, which in turn triggers the first sub-transaction, labeled ”Start Sub-Transaction: Verify Funds.” The purpose of this initial sub-process is to confirm the availability of adequate funds through a ”CheckBalance” operation. Based on the outcome of this check, the system either proceeds to a ”SufficientFunds” state, culminating in a successful ”Commit Sub-Transaction: Verify Funds,” or encounters ”InsufficientFunds,” leading to an ”Abort Sub-Transaction: Verify Funds.”. The subsequent step, ”Start Sub-Transaction: Transfer Funds,” is initiated solely upon the successful completion (commit) of the ”Verify Funds” sub-transaction. This second sub-process involves the core actions of ”Debit Account A” and ”Credit Account B.” A successful execution of both these actions results in a ”Commit Sub-Transaction: Transfer Funds.” Conversely, any failure during either the debit or credit operation leads to an ”Abort Sub-Transaction: Transfer Funds.”. The success of the entire transaction, spanning from the initial ”Start Transaction” to the final ”Commit Transaction,” is contingent upon the successful commitment of both constituent sub-transactions. If either the ”Abort Sub-Transaction: Verify Funds” or the ”Abort Sub-Transaction: Transfer Funds” occurs, the entire operation is rolled back, resulting in an ”Abort Transaction.” This mechanism ensures the integrity of the data by reverting any partial modifications. Consequently, the final status of the primary transaction is directly determined by the combined outcomes of its sub-transactions.
Figure 4.8: Nested Transaction containing sub-transactions depicting a bank transaction.
@@ 4.6.1 Atomic Commit Protocols
Two-Phase Commit (2PC) protocol permits any party (participant) can decide to roll back its transaction (due to failure) and the entire transaction is globally reverted [coulouris2012, tanenbaum2007]. This requires a coordinator to facilitate concurrency control. The algorithm has two phases [coulouris2012]: in the first phase, each party votes to commit or abort a transaction. Once, a party has voted to commit a transaction, it is binding and immutable. After voting each party goes to the ”prepared” state. In the second phase, each party in the transaction jointly executes the decision. If any party aborts or fails, then the overall transaction is aborted. The two phases ensure that all parties reach the same decision on committing or rolling back the transaction.
@@
Figure 4.9: Two-Phase Commit [coulouris2012].
@@
Two-Phase Commit fails when the coordinator is dead and participants in the transaction cannot determine the state of the transaction resulting in longer polling intervals as parties try to connect and get information from the coordinator. Three-Phase Commit solves the problem by taking more rounds [coulouris2012, tanenbaum2007] as we perform a precommit round and obtain ACKS before proceeding to commit.
@@
Figure 4.10: Three-Phase Commit [coulouris2012].
@@ 5.1 Gossip Algorithm
The Gossip protocol [Ganesh2003, Jelasity2007, Kermarrec2007, Leito2010] presents a robust and widely applicable anti-entropy mechanism that can be effectively employed in distributed systems to promote data consistency [petrov2019]. This protocol leverages a probabilistic approach to disseminate information (messages) across a network, drawing a compelling analogy to the way rumors or information propagate through a social network. A significant advantage of the Gossip algorithm is its inherent fault tolerance and resilience to failures, even in the face of substantial and unpredictable network disruptions or node outages. While the epidemic approach, characterized by the rapid and potentially ubiquitous spread of information to all reachable nodes, is a common and intuitive implementation strategy, Gossip protocols can also be implemented using non-epidemic strategies that strategically leverage overlay network structures or maintain partial views of the system’s membership to optimize message dissemination and resource utilization.
@@ 5.2 CRDT
Conflict-free Replicated Data Types (CRDTs) represent another powerful and increasingly popular class of anti-entropy mechanisms in distributed systems. Riak, a distributed NoSQL database, was a notable pioneer in the widespread adoption and practical application of CRDTs [crdt2022, petrov2019]. A CRDT is a specialized type of data structure meticulously designed to maintain a consistent state across multiple independent replicas, regardless of the order in which concurrent operations are executed on those individual replicas. This inherent order-independence property makes CRDTs exceptionally well-suited for scenarios involving collaborative applications and seamless remote synchronization (replication) of data across numerous geographically distributed devices, a task that can be inherently complex to manage for data consistency using traditional concurrency control mechanisms. Mathematically, CRDTs can be formally characterized as partially ordered monoids that possess a well-defined lattice structure. To ensure their conflict-free nature, they must satisfy the algebraic properties of commutativity (the order in which operations are applied does not affect the final state) and idempotence (applying the same operation multiple times has the same effect as applying it only once).
@@
A diverse range of CRDTs has been developed, each tailored to manage different types of data and support specific sets of operations. Common examples include map CRDTs (such as the Last-Write-Wins Map - LWW-Map), set CRDTs (such as the Grow-Only Set - G-Set), and many other specialized data structures designed for various consistency and concurrency control requirements [crdt2022, petrov2019].
@@
The provided C source code implements a Grow-only Counter (GCounter) as shown in crdt-gcounter.c, a type of Conflict-free Replicated Data Type (CRDT), utilizing the Message Passing Interface (MPI) for parallel operation. Each parallel process maintains a local state stored in the ‘GCounter‘ structure, where the ‘counts‘ array tracks the contributions of every process on a per-index level. The core CRDT operation is the merge, which is performed using the collective primitive ‘MPI_Allreduce‘ with the reduction operator ‘MPI_MAX‘. This operation ensures that every process receives the global state where each element of the merged array is the maximum of the corresponding local element from all contributing arrays, correctly satisfying the GCounter’s merge property: $merge(A,B)=max(A_{i},B_{i})$. Finally, the ‘gcounter_value‘ function calculates the counter’s total by performing a summation over all indices of the merged array; while the true CRDT merge logic uses ‘MPI_MAX‘, the summation provides a shorthand method to check that the merged arrays have similar content and allows for a final, consistent global value check across all processes for demonstration purposes.
@@ 5.5.2 Multiple Description Coding
 [^1] See an illustrative example of optimizing a distributed data store for operation at significant scale in the Uber blog
@@
Peer-to-peer computing offers notable benefits, including reduced storage and computational costs, achieved by distributing the burden of computation and data storage across several participating nodes. The applications of peer-to-peer computing span a wide range of domains, including distributed file systems (e.g., OceanStore [Kubiatowicz2000], PAST [Druschel2001], CFS [DabekKKMS2001]), file sharing networks (e.g., BitTorrent [BitTorrent2019], FastTrack [FastTrack2022]), and large-scale Grid computing platforms [GridComputing2022].
@@ 6.1 Resilient Overlay Networks
An overlay network provides a dynamic, just-in-time network infrastructure built on top of an existing network layer (like the Internet) to facilitate decentralized communication in a scalable manner. It achieves this by implementing its own routing and packet transfer functionalities at the application level, abstracting away the complexities of the underlying physical network [caesar2006, david2001].
Overlay networks must be designed to handle a high rate of client disconnections, which can occur due to various factors such as poor network conditions, limited available bandwidth, and the need for dynamic balancing of processing and bandwidth capacity across the participating nodes [castro2005]. Consequently, there is a significant need for the development of resilient overlay networks that enhance the reliability of the network and make it robust to node and link failures [david2001]. Nodes within resilient overlay networks often incorporate self-healing capabilities. Such nodes can perform quality of service (QoS) checks on the communication links they use and dynamically re-route packets as necessary to circumvent failures and reduce latency.
@@
When communicating across different Internet Service Providers (ISPs), one might naively assume that Border Gateway Protocol (BGP) routing records are consistently up-to-date and reflect the optimal paths. However, this is often not the case, as the entries in an ISP’s routing table can be influenced by a multitude of factors, including financial agreements between ISPs, the time it takes for BGP to converge after a routing update, and even political considerations [david2001].
The routing table maintained at the application level by the overlay network can potentially identify and utilize more efficient and reliable communication links compared to the underlying BGP routes. Content Delivery Networks (CDNs) are a prominent example of how overlay networks are leveraged to improve offsite caching and the efficient delivery of web resources to end-users [david2001].
@@ 6.2 Unstructured P2P Systems
Unstructured overlay networks are characterized by the absence of a predefined, deterministic structure for organizing nodes and data. Consequently, they do not inherently provide efficient query-matching mechanisms for locating specific data within the network [castro2005]. Examples of P2P systems that utilize unstructured overlays with a random graph topology include Gnutella [taylor2009], the now-defunct Kazaa, and Freenet [clarke2001]. In such networks, a typical search query is flooded across the network to a subset or all neighboring nodes. Each node receiving the query then compares it against the metadata or content it holds.
@@ 6.3 Structured P2P Systems
In contrast to unstructured P2P systems, structured overlay networks impose a specific, well-defined structure on the network topology to facilitate efficient query matching and data discovery [castro2005]. This efficient search capability is achieved by enforcing constraints on how the network is constructed and how data is indexed and located. In a structured P2P system, each node is typically associated with a unique key or identifier, and the actual data or metadata is stored at the node whose key is closest to the key of the data item. Searching for a specific data item involves querying the network using its associated key. The network’s underlying structure, often based on sophisticated data structures like Distributed Hash Tables (DHTs), ensures that queries can be routed efficiently to the node(s) likely to hold the desired information. Prominent examples of P2P systems that leverage structured overlay networks include Chord [stoica2003], Tapestry [zhao2004], Kademlia [kademlia], and Pastry [rowstron2001].